1. A boy stands on a bridge holding an apple. He eats the apple, then throws the core up in the air at a velocity of 15 m/s. Set t=0 to to the instant after the apple leaves his hand, and take acceleration during free fall to be -10 m/s. Complete the following table: time velocity acceleration (seconds) (m/s) (m/s2) t=0 -15 m/s t = 3.7 In the space below, draw complete graphs of position vs. time, velocity vs. time, and acceleration vs. time for the first 3.7 seconds of free fall. You must label your vertical and horizontal axes on all three graphs, including the units. Also label any r- and y-intercepts. Take the boys hand to be position zero.
Answer #1

At t=0, v= 15m/s and a= -10 m/s2

At t=1 s,

v= u+at= 15 -10= 5m/s and a= -10m/s2

When velocity= -15m/s,

v= u+at

So, -15= 15 - 10t

t= 3s and a= -10m/s2

At t= 3.7 s,

v= u+at.

v= 15- 37= -22 m/s and a= -10m/s2.

11.5 Displacement 10 1.5 Time (s) And at 3.7s, graph goes downward to a displacement of -12.95mVelocity (m/s) 15 3.7 Time (s) C) 1.5 -15 -22Acceleration (m/s) Time (s) 10